When the Crampton coal-fired train engine was built in 1852, it's mass was 48.3 tonnes and its force capability was rated at 22.4 kN. Assuming it was pulling train cars whose total mass doubled its own mass and the total friction on the engine and cars was 10.1 kN,what was the magnitude of the train's acceleration?
1
Expert's answer
2020-04-23T12:54:10-0400
"2ma=F-F_f"

"2(48300)a=22400-10100"

"a=0.127\frac{m}{s^2}"